OpenStudy (anonymous):

Add the mixed numbers 3 2/3 and 2 4/7 by converting each into an improper fraction.

OpenStudy (anonymous):

please help me with this

OpenStudy (anonymous):

As improper fractions, 3 2/3 = 11/3 and 2 4/7 = 18/7 In order to add them together, you also need to have a common denominator, which would be 21. 3 2/3 = 11/3 = 77/21 and 2 4/7 = 18/7 = 54/21 77/21 + 54/21 = 131/21 131/21 = 6 5/21

OpenStudy (anonymous):

Can you suggest a alternative way to do this problem

OpenStudy (anonymous):

and how did you get 131/21 to 6 5/21